# Cascade-Zero123: One Image to Highly Consistent 3D with Self-Prompted Nearby Views

### v1

@@ 3.5 Self-Distillation Design
Considering that the Refiner-0123 network $f_{\phi_{r}}$ is trained via back-propagation to minimize the denoising loss. Inspired by the methods [6, 16, 3, 7] in the self-supervised learning field. The Base-0123 network is updated in a momentum update way using exponential moving average (EMA). Specifically, we have denoted the parameters of $f_{\phi_{b}}$ as $\phi_{b}$ and those of $f_{\phi_{r}}$ as $\phi_{r}$ we update $\phi_{b}$ by
@@ 3.6 Inference
During the inference stage, given an input image, we need to go through two stages of the DDIM (Denoising Diffusion Probabilistic Model) [19] schedule while simultaneously passing the input image through both the Base-0123 and Refiner-0123 networks.
@@ 4 Experiments
Figure 3: Novel view synthesis compared with Zero123-XL [9], and SyncDreamer [31], where Zero123-XL is Zero-1-to-3 pre-trained on Objaverse-XL datasets [9], achieving higher performance. We selected some challenging scenes, including stacked objects, parallel objects, and objects with multiple branches. Zero-1-to-3 exhibits good quality in image generation but lacks consistency in these complex scenes. SyncDreamer, on the other hand, demonstrates good consistency but struggles to maintain good quality in image generation. Our model, however, maintains both quality and consistency in these scenarios.
We assess our model’s performance on novel view synthesis and single-image to 3D reconstruction tasks. We introduce the datasets, implementation details and metrics from Sec.4.1 to Sec.4.3. And show qualitative results, quantitative results, and ablation studies from Sec.4.4 to Sec.4.6
@@ 4.1 Datasets
Objaverse Dataset is a large-scale dataset containing 800K+ annotated 3D mesh objects [39]. We use this dataset for training and validation. We render 12 images per object from uniformly distributed viewpoints surrounding the object, followed by Zero-1-to-3.
Realfusion15 Realfusion15 is the dataset collected and released by RealFusion [34], consisting of 15 natural images that include bananas, birds, cacti, barbie cakes, cat statues, teapots, microphones, dragon statues, fishes, cherries, and watercolor paintings, etc.
@@ 4.2 Implementation Details
We train Cascade-Zero123 on the Objaverse dataset [10] which contains about 800k objects. Following Zero-1-to-3, the number of viewpoints rendered is 12. Following the assumption of Zero-1-to-3, we also assume that the azimuth of both the input view and the first target view is $0^{\circ}$. We train the Cascade-Zero123 for 170k steps with 8 V100 GPUs using a total batch size of 96.
@@ 4.3 Metrics
We use the following evaluation metrics to quantitatively evaluate the performance of our model. We first report the Peak Signal-to-Noise Ratio (PSNR). Perceptual Loss (LPIPS) measures the perceptual distance between two images by comparing the deep features extracted by deep neural networks given each image as input. Structural Similarity (SSIM) measures the structural similarity between two images considering both color and texture information. The CLIP-score quantifies the average CLIP distance between the rendered image and the reference image, serving as a measure of 3D consistency by assessing appearance similarity across novel views and the reference view.
Figure 4: Single image to 3D reconstruction using SDS loss [42] compared with Zero123-XL. The first two rows illustrate that Cascade-Zero123 can correct the problem that the Zero-1-to-3 model sometimes learns inaccurate colors of the backside. The middle two lines describe how Cascade-Zero123 can rectify structural errors in the Zero-1-to-3 model through multi-view self-prompting. The last two lines indicate that Cascade-Zero123 can address the problem of transparent or high-brightness objects being mistakenly learned as white clouds.
@@ 4.4 Qualitative Results
Single Image to 3D Using SDS Loss We show qualitative results of a single image to 3D using Dreamfusion’s Score Distillation Sampling (SDS) loss [42]. All samples are generated by our Cascade-Zero123 and baseline Zero123-XL in Fig. 4. Firstly, as shown in the upper two rows of Fig. 4, after using SDS loss, Zero-1-to-3 tends to learn the backside as the smooth color of the front side. This is due to the one-to-one sparsely generation pipeline so that the model can only perceive information from the prompt of a single image. As a result, the ears of the fox, the back and hair of Lisa(the woman in yellow), and the back of the dinosaur are all learned as the same color. Additionally, due to the influence of front-facing lighting, the top of the little Batman’s head is learned as pure white. However, it should be black, the same as the back. With the correction provided by self-prompted views, we no longer have to speculate the color of the backside from scratch. Instead, we can progressively predict the backside based on the side views. This allows us to better learn the true color consistency of unseen views.
In addition, the proposal of self-prompted views has enhanced the ability of Zero-1-to-3 to model the shapes of multiple objects and occluded objects as shown in the middle of Fig. 4. The shapes of multiple bottles are well preserved, and even the occluded rocket fins can be modeled. Moreover, Zero-1-to-3, when combined with SDS, has difficulty modeling objects with transparency. Zero-1-to-3 sometimes learns the backside of transparent or high-brightness objects as pure white mist-like clouds. However, the multi-view conditioning information prevents the model from lacking information about other sides, avoiding the learning of a pure white plane for the backside and providing 3D information for objects with transparency.
@@
Methods $\uparrow$ PSNR $\uparrow$ SSIM $\downarrow$ LPIPS $\uparrow$ CLIP-Score
@@
Methods $\downarrow$ LPIPS $\uparrow$ CLIP-Score
RealFusion [34] 0.197 0.735
Make-it-3D [56] 0.119 0.839
Zero-1-to-3 [29, 9] 0.068 0.759
Magic123 [43] 0.062 0.747
Consistent-123 [27] 0.056 0.844
Ours 0.043 0.916
@@ 4.5 Quantitative Results
Following prior research [60], which randomly picked up 100 objects from the Objaverse testset, Since the entire Objaverse test set [10] is quite large, testing all samples would require an excessively long time. Following the approach of Consistent123 [60], we randomly selected a subset of samples from the Objaverse test set for testing. These selected samples have not been seen during the training process. Compared with Consistent123 [60], we take a larger number of samples for evaluation (Consistent123 only takes 100 samples). Specifically, we randomly selected 200 samples and then randomly selected an input view and a target view following the same setting as Zero-1-to-3. On the Objaverse test set, we measured the reconstruction performance and consistency using metrics of PSNR, SSIM, LPIPS, and CLIP-score. We utilize the checkpoints from Zero123-XL [29, 9]. The index of the selected samples will be made available in the open-source release.
@@
Following Magic123 [43] and Consistent123 [27] We evaluate Cascade-Zero123 against state-of-the-art baselines, including RealFusion [34], Make-it-3D [56], Zero-1-to-3 [29] and Magic123 [43], on the RealFusion15 datasets. Like Magic123, we measure the CLIP-scores and LPIPS which assess the reconstruction quality and visual consistency. As shown in Table. 2, our Cascade-Zero123 surpasses the of state-of-the-art baselines by a lot.
Table 3: Ablation studies on the modules of Cascade-Zero123. We evaluate our method on the test split of Objaverse [10].
$\uparrow$ PSNR $\downarrow$ LPIPS $\uparrow$ CLIP-Score
Zero123-XL 18.68 0.189 0.758
+Cascade two Zero-1-to-3 (frozen) 17.67 0.189 0.778
 +Prompted random views 18.55 0.179 0.787
 +Prompted larger views 19.04 0.172 0.783
 +Prompted more nearby views 19.13 0.165 0.789
 +Prompted nearby views (Ours) 20.04 0.137 0.800
   -coarse target views 19.99 0.143 0.792
   -EMA to Base-0123 19.89 0.139 0.801
4.6 Ablation Studies
@@
As shown in Table 3, in the ablation study, we provide a detailed explanation for the effectiveness of our method, Cascade-Zero123, based on the following aspects. Firstly we will study whether nearby views perform better than randomly selected views or larger rotated views. Then we will study the impact of having larger rotated prompted condition views. We will also explore the effects of not using the Exponential Moving Average (EMA) and coarse target views. All the experiments presented below were conducted using the same model, pretraining for 120k iterations on 800K samples from the Objaverse dataset. We experiment with the same testing protocol on the Objaverse testset.
@@
First, we will validate the difference between incorporating prompted views as inputs to the second-stage model compared to using only the fully trained first-stage model. This analysis will measure the gain achieved by the design of Cascade-Zero123 over the standalone Zero123-XL model. We will then explore various ablations of self-prompted views, including completely random prompted views, denoted as ”prompted random views”; inputting images from surrounding viewpoints, specifically inputting images generated by the Base-0123 at $[90^{\circ},180^{\circ},270^{\circ}]$ angles as the condition images for the second stage, denoted as ”prompted larger views”; and the effects of having more prompted views on the second-stage model. In practice, we selected the azimuth angles of $[-135^{\circ},-90^{\circ},-45^{\circ},45^{\circ},90^{\circ},135^{\circ}]$ and elevation angles of $[-10^{\circ},30^{\circ}]$. We also conduct ablation on the setting of without target views being part of the self-prompted views. Our model results are highlighted in blue.
By simply cascading two Zero-1-to-3 models without training and using self-prompted nearby views, we were able to improve the LPIPS and CLIP scores to some extent, but the PSNR decreased. This validates our hypothesis that using multiple viewpoints as conditions can enhance consistency, but simply cascading them may not work. Secondly, using the same number of random views, larger view degree views or even more views with larger view degree views somewhat do harm to the performance, indicating the need for Zero-1-to-3 to progressively rotate the camera views. Making a large rotation in views at once would lead to large performance degradation.
@@
When the parameters of Base-0123 are not updated and kept fixed, the performance will also degrade. The design of EMA is also shown to have a certain positive effect.
5 Limitation and Conclusion
Limitation
Cascade-Zero123 is based on the pre-trained Zero-1-to-3 model [29]. For cases that are extremely difficult for Zero-1-to-3, Cascade-Zero123 has limited ability to handle. The first one lies in the deeply occluded samples. With 2D image input, it is challenging to figure out the exact overlap, so even with nearby views, the overlapping parts may still appear to be stuck together. The 3D depth information will be lost and a flat structure will be potentially learned. Additionally, Zero-1-to-3 is sensitive to camera pose elevation. Therefore, Cascade-Zero123 also struggles with input images that have a high elevation. While our model has already achieved performance and consistency improvement compared to previous models, there are still some potential future directions, e.g. enhancing Cascade-Zero123 by incorporating attention between multiple views as [54] and involving multi-modal conditioning information, e.g. depth or normal.
Conclusion
In this paper, we propose a Cascade-Zero123 network using self-prompted nearby views. Cascade-Zero123 utilizes Zero-1-to-3 as an additional prior constraint, which modifies the one-to-one generation mode of Zero-1-to-3 and retains the generalized capability as well. With the proposed self-prompted mode, self-prompted views are used as input without the need for additional modules for training. This allows Cascade-Zero123 to improve the geometric and visual consistency of Zero-1-to-3.
Appendix A Appendix
A.1 More Qualitative Results of Single Image to 3D
@@
We show more qualitative results of the single image to 3D using Dreamfusion’s Score Distillation Sampling (SDS) loss [42]. All samples are generated by our Cascade-Zero123 and baseline Zero123-XL in Fig. 5. As shown in the figure, Cascade-Zero123 is able to correct structural and textural errors, as well as the backside errors of objects. For example, the first row of the figure shows that Cascade-Zero123 fixes the texture and structure errors on the backside of the ice cream. And Cascade-Zero123 correct the color of the backside of the pears and the structure of the pear pedicle at the same time. With the correction provided by self-prompted views, we no longer have to speculate the color of the backside from scratch. Instead, we can progressively predict the backside based on the side views. This allows us to better learn the true color consistency of unseen views.
@@
We also provide qualitative results of multi-view conditioning information that can prevent the incorrect model of transparent objects or objects with high brightness. Our methods avoid learning a pure white plane for the backside and provide more accurate 3D information for these objects. For example, the butterfly wings and the icicle have been learned as solid blocks instead.
A.2 More Qualitative Results of Novel View Synthesis

### v2

@@ 3.5 Self-Distillation Design
Considering that the Refiner-0123 network $f_{\phi_{r}}$ is trained via back-propagation to minimize the denoising loss. Inspired by the methods [9, 23, 5, 11] in the self-supervised learning field. Our ablation experiments demonstrated that even if only accepted a single image, Refiner-0123 still improves performance. To further enhance the Base-0123 model, the Base-0123 network is updated in a momentum update way using exponential moving average (EMA). Specifically, we have denoted the parameters of $f_{\phi_{b}}$ as $\phi_{b}$ and those of $f_{\phi_{r}}$ as $\phi_{r}$ we update $\phi_{b}$ by
@@ 3.6 Inference
During the inference stage, given an input image, two stages of the DDIM (Denoising Diffusion Probabilistic Model) [28] schedule are performed. The input image passes through both the Base-0123 and Refiner-0123 simultaneously.
@@ 4 Experiments
Figure 4: Novel view synthesis compared with Zero123-XL [15], and SyncDreamer [46], where Zero123-XL is Zero-1-to-3 pre-trained on Objaverse-XL datasets [15], achieving higher performance. We selected some challenging scenes, including stacked objects, parallel objects, and objects with multiple branches. Zero123-XL exhibits good quality in image generation but lacks consistency in these complex scenes. SyncDreamer demonstrates good consistency but struggles to maintain good quality in image generation. Our model, however, maintains both quality and consistency in these scenarios.
We assess our model’s performance on novel view synthesis and single-image to 3D reconstruction tasks. We introduce the datasets, implementation details, and metrics from Sec.4.1 to Sec.4.3. And show qualitative results, quantitative results on Sec.4.4. Ablation studies are introduced in supplementary materials.
@@ 4.1 Datasets
Objaverse Dataset is a large-scale dataset containing 800K+ annotated 3D mesh objects [55]. We use this dataset for training and validation. We render 12 images per object from uniformly distributed viewpoints followed by [43].
Realfusion15 Realfusion15 is the dataset collected and released by RealFusion [50], consisting of 15 natural images.
Google Scanned Object (GSO) [18] is a high-quality scanned household items dataset. We adopt the same data split as in SyncDreamer [46].
RTMV [81] consists of complex scenes, and each scene is composed of 20 random objects. We adopt the same data split as in Zero-1-to-3 [43].
@@ 4.2 Implementation Details
We train Cascade-Zero123 on the Objaverse dataset [16] which contains about 800k objects. Following Zero-1-to-3, the number of viewpoints rendered is 12. Following the assumption of Zero-1-to-3, we also assume that the azimuth of both the input view and the first target view is $0^{\circ}$. We train the Cascade-Zero123 for 200k steps with 8 V100 GPUs using a total batch size of 96.
@@ 4.3 Metrics
We use the following evaluation metrics to quantitatively evaluate the performance of our model. We first report the Peak Signal-to-Noise Ratio (PSNR). Perceptual Loss (LPIPS) measures the perceptual distance between two images by comparing the deep features extracted by deep neural networks given each image as input. Structural Similarity (SSIM) measures the structural similarity between two images considering both color and texture information. The CLIP-score (CLIP) quantifies the average CLIP distance between the rendered image and the reference image, serving as a measure of 3D consistency by assessing appearance similarity across novel views and the reference view. We measured Chamfer Distance (C Dist.) to evaluate point-by-point shape similarity and volumetric IoU (Vol IoU) to quantify the overlap between the reconstructed and the ground truth shape.
Figure 5: Single image to 3D reconstruction using SDS loss [60] compared with Zero123-XL. The first two rows illustrate that Cascade-Zero123 can correct the problem that Zero-1-to-3 sometimes learns inaccurate colors of the backside. The middle two lines describe how Cascade-Zero123 can rectify structural errors through multi-view self-prompting. The last two lines indicate that Cascade-Zero123 can address the problem of transparent or high-brightness objects being mistakenly learned as white clouds.
@@ 4.4 Qualitative Results
Single Image to 3D Using SDS Loss We show qualitative results of a single image to 3D using Dreamfusion’s Score Distillation Sampling (SDS) loss [60]. All samples are generated by our Cascade-Zero123 and baseline Zero123-XL in Fig. 5. Firstly, as shown in the upper two rows of Fig. 5, after using SDS loss, Zero-1-to-3 tends to learn the backside as the smooth color of the front side. This is due to that the model can only perceive information from the prompt of a single image. As a result, the ears of the fox, the back and hair of Lisa (the woman in yellow), the back of the dinosaur, and the glass tiger are all learned as the same color. With the correction provided by self-prompted views, we no longer have to speculate the color of the backside from scratch. Instead, we can progressively predict the backside based on the side views. This allows us to better learn the true color consistency of unseen views.
In addition, the proposal of self-prompted views has enhanced the ability of Zero-1-to-3 to model the shapes and textures of objects as shown in the middle of Fig. 5. The shapes of the ice cream are well preserved, and even the complex structure of the ninja’s back can be modeled. Moreover, Zero-1-to-3, when combined with SDS, has difficulty modeling objects with transparency. Zero-1-to-3 sometimes learns the backside of transparent or high-brightness objects as pure white mist-like clouds. However, the multi-view conditioning information prevents the model from lacking information about other sides, avoiding the learning of a pure white plane for the backside and providing 3D information for objects with transparency.
@@
[…]
@@
Table 3: Evaluation of novel view synthesis on GSO dataset [18].
[…]
Realfusion[50] 15.26 0.722 0.283 -
Zero-1-to-3[43] 18.93 0.779 0.166 -
SyncDreamer[46] 20.05 0.798 0.146 -
Zero123-XL[15] 18.81 0.828 0.142 0.813
Ours(Zero123-XL) 20.35 0.850 0.113 0.846
Ours(SyncDreamer) 21.10 0.862 0.095 0.944
Table 4: Evaluation of novel view synthesis on RTMV dataset [81].
[…]
DietNeRF [27] 7.13 0.406 0.507 -
Image Variation [1] 6.56 0.442 0.564 -
SJC-I [85] 7.95 0.456 0.545 -
Zero123-XL[15] 10.97 0.573 0.465 0.702
Ours 11.12 0.617 0.441 0.734
@@
[…]
RealFusion [50] 20.216 0.197 0.735
Make-it-3D [79] 20.010 0.119 0.839
Zero-1-to-3 [43] 25.386 0.068 0.759
Zero123-XL [15] 25.220 0.050 0.897
Magic123 [61] 25.637 0.062 0.747
Consistent-123 [40] 25.682 0.056 0.844
Ours 26.098 0.043 0.916
Table 6: Evaluation of single image to 3D reconstruction on GSO dataset [18].
[…]
Realfusion [50] 0.0819 0.2741
Magic123 [61] 0.0516 0.4528
One-2-3-45 [42] 0.0629 0.4086
Point-E [54] 0.0426 0.2875
Shap-E [30] 0.0436 0.3584
Zero123 [43] 0.0339 0.5035
SyncDreamer [46] 0.0261 0.5421
Ours 0.0207 0.5792
@@ 4.5 Quantitative Results
Novel view synthesis on Objaverse testset. Following prior research [88], which randomly picked up 100 objects from the Objaverse testset, Since the entire Objaverse test set [16] is quite large, testing all samples would require an excessively long time. Following the approach of Consistent123 [88], we randomly selected a subset of samples from the Objaverse test set for testing. These selected samples have not been seen during the training process. Compared with Consistent123 [88], we take a larger number of samples for evaluation (Consistent123 only takes 100 samples). Specifically, we randomly selected 200 samples and then randomly selected an input view and a target view following the same setting as Zero-1-to-3. On the Objaverse test set, we measured the reconstruction performance and consistency using metrics of PSNR, SSIM, LPIPS, and CLIP-score (CLIP). We utilize the checkpoints from Zero123-XL [43, 15]. The index of the selected samples will be made available in the open-source release.
Novel view synthesis on Google Scanned Object (GSO). Following SyncDreamer [46] and Zero-1-to-3 [43], we adopt the Google Scanned Object [18] dataset as the evaluation dataset. We followed the setting of SyncDreamer [46] that chose the same 30 objects ranging from daily objects to animals of Google Scanned Object. We also apply the proposed cascade mechanism to SyncDreamer [46], named as Ours(SyncDreamer). The self-prompted views generated from the first-stage Base-SyncDreamer are fed into the Synchronized Multiview Noise Predictor in the second-stage Refiner-SyncDreamer, as shown in Table 4. We replicate Zero123-XL [15] for fair comparison. We measured the reconstruction performance and consistency using metrics of PSNR, SSIM, LPIPS, and CLIP-score (CLIP). Our cascade method also shows notable improvement on both Zero123-XL and SyncDreamer.
Novel view synthesis on RTMV. In Table 4, we use the same experimental setup as Zero1-to-3 [43] on RTMV [81] and replicate the performance of the Zero123-XL [15]. We measured the reconstruction performance and consistency using metrics of PSNR, SSIM, LPIPS, and CLIP-score (CLIP).
@@
Novel view synthesis on Realfusion15 testset. Following Magic123 [61] and Consistent123 [40] We evaluate Cascade-Zero123 against many related baselines, including RealFusion [50], Make-it-3D [79], Zero-1-to-3 [43] and Magic123 [61], on the RealFusion15 datasets. Like Magic123, we measure the PSNR, LPIPS, and CLIP-scores (CLIP), which assess the reconstruction quality and visual consistency. As shown in Table. LABEL:quantitative_realfusion15.
Single image to 3D reconstruction on GSO. In Table 6, we compare our approach with various other methods following the setting of SyncDreamer [46]. Both our method and SDS-free methods utilize NeuS [86], a neural reconstruction method for converting multi-view images into 3D shapes. To achieve faithful reconstruction of 3D mesh that aligns well with ground truth, the generated multi-view images should be geometrically coherent. Cascade-Zero123 achieves the best results in both Chamfer distance (C Dist.) and volumetric IoU (Vol IoU) metrics, demonstrating proficiency in producing multi-view consistent images.
5 Limitation and Conclusion
Limitation
Cascade-Zero123 is based on the pre-trained Zero-1-to-3 model [43]. For cases that are extremely difficult for Zero-1-to-3, Cascade-Zero123 has limited ability to handle. With 2D image input, it is challenging to figure out the exact overlap, so even with nearby views, the overlapping parts may still appear to be stuck together. The 3D depth information will be lost and a flat structure will be potentially learned. Additionally, Zero-1-to-3 is sensitive to camera pose elevation. Therefore, Cascade-Zero123 also struggles with input images that have a high elevation. While our model has already achieved performance and consistency improvement compared to previous models, there are still some potential future directions, e.g. enhancing Cascade-Zero123 by incorporating attention between multiple views as [72] and involving multi-modal conditioning information, e.g. depth or normal.
Conclusion
We have witnessed Zero-1-to-3’s success in synthesizing novel views from a single image. However, inconsistency still appears as the input information is too limited. We propose a Cascade-Zero123 network to progressively dig out the multi-view information from the input image by self-prompting nearby views. A series of experimental proofs have demonstrated the effectiveness of this cascade design. We believe the cascade framework can be compatible with more Zero-1-to-3 variants and leave this as an important future work.
Acknowledgements
This work was supported in part by the National Natural Science Foundation of China under Grant 62125109, Grant 61931023, Grant 61932022, Grant 62371288, Grant 62320106003, Grant 62301299, Grant T2122024, Grant 62120106007.
Appendix 0.A Appendix
0.A.1 Ablation Studies
@@
As shown in Table LABEL:tab:ablation, in the ablation study, we provide a detailed explanation for the effectiveness of our method, Cascade-Zero123, based on the following aspects. Firstly we will study whether nearby views perform better than randomly selected views or larger rotated views. Then we will study the impact of having larger rotated prompted condition views. We will also explore the effects of not using the Exponential Moving Average (EMA) and coarse target views. All the experiments presented below were conducted using the same model, pretraining for 12k iterations on 800K samples from the Objaverse dataset. We experiment with the same testing protocol on the Objaverse testset.
@@
First, we will validate the difference between incorporating prompted views as inputs to the second-stage model compared to using only the fully trained first-stage model. This analysis will measure the gain achieved by the design of Cascade-Zero123 over the standalone Zero123-XL model. We will then explore various ablations of self-prompted views, including completely random prompted views, denoted as "prompted random views"; inputting images from surrounding viewpoints, specifically inputting images generated by the Base-0123 at $[90^{\circ},180^{\circ},270^{\circ}]$ angles as the condition images for the second stage, denoted as "prompted larger views"; and the effects of having more prompted views on the second-stage model. In practice, we selected the azimuth angles of $[-135^{\circ},-90^{\circ},-45^{\circ},45^{\circ},90^{\circ},135^{\circ}]$ and elevation angles of $[-10^{\circ},30^{\circ}]$. We also conduct ablation on the setting without target views being part of the self-prompted views. Our model results are highlighted in blue.
By simply cascading two Zero-1-to-3 models without training and using self-prompted nearby views, we were able to improve the LPIPS and CLIP scores to some extent, but the PSNR decreased. This validates our hypothesis that using multiple viewpoints as conditions can enhance consistency, but simply cascading them may not work. Secondly, using the same number of random views, larger view degree views or even more views with larger view degree views somewhat harm the performance, indicating the need for Zero-1-to-3 to progressively rotate the camera views. Making a large rotation in views at once would lead to large performance degradation.
@@
When the parameters of Base-0123 are not updated and kept fixed, the performance will also degrade. The design of EMA is also shown to have a certain positive effect. Furthermore, As shown in “Only single-image prompt for Refiner-0123”, even if Refiner-0123 accepts a single image, it still outperforms Zero123-XL. So Base-0123 will benefit by inheriting parameters from Refiner-0123.
0.A.2 More Qualitative Results of Single Image to 3D
@@
We show more qualitative results of single image to 3D Dreamfusion’s Score Distillation Sampling (SDS) loss [60]. All samples are generated by our Cascade-Zero123 and baseline Zero123-XL in Fig. 6. As shown in the figure, Cascade-Zero123 is able to correct structural and textural errors, as well as the backside errors of objects. For example, the first row of the figure shows that Cascade-Zero123 fixes the texture and structure errors on the backside of the ice cream. And Cascade-Zero123 correct the color of the backside of the pears and the structure of the pear pedicle at the same time. With the correction provided by self-prompted views, we no longer have to speculate the color of the backside from scratch. Instead, we can progressively predict the backside based on the side views. This allows us to better learn the true color consistency of unseen views.
@@
We also provide qualitative results of multi-view conditioning information to prevent the model from lacking information about transparency, avoiding the learning of a pure white plane for the backside and providing 3D information for objects with transparency. For example, the butterfly wings have been learned as solid white blocks instead.
0.A.3 More Qualitative Results of Novel View Synthesis